OpenStudy (chmvijay):

I Would like to recommend @somy for the Ambassador as she is doing very well in chemistry blog :) hope they will soon :) Thank you ...

OpenStudy (abhisar):

Yes me too. @Somy is doing very hard work in chemistry. She is very talented in biology too.

OpenStudy (somy):

OMG Umm thank you guys a lot!!!!! but umm i don't think i can do it <.<

OpenStudy (anonymous):

Somy did you apply, if not you should :D

OpenStudy (shrutipande9):

one more vote for @somy i recommend @abhisar too

OpenStudy (anonymous):

^^^^^^^

OpenStudy (beccaboo333):

This isn't that much different than asking for someone to become a moderator. Somy still needs to fill out an application to become an ambassador. You can give the request to the Ambassador Team Moderators. As for soon to be ambassadors, they must be able to fill the parts of what's needed to become an ambassador. I do too agree that somy could be a great ambassador, but there is a lot to even become one. @somy check out the ambassador assignments we had a while back located in the ambassador section. It may help you decide if you'd truly like to try to become one. The Ambie Team leaders are @hartnn @e.mccormick @blues @Kayne @thomaster

OpenStudy (abhisar):

Thank you @shrutipande9 ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (compassionate):

I vote @somy for System Administrator.

OpenStudy (e.mccormick):

We do not just add people to the Ambassador program. We have an application and training process.

OpenStudy (beccaboo333):

Like what I had said @e.mccormick ._.

OpenStudy (somy):

umm @chmvijay @Abhisar @shrutipande9 @iambatman @Compassionate guys Thanks a lot!!! im sorry but i don't think i will do it T_T

OpenStudy (goformit100):

My vote too goes for Sir @Abhisar And Ma'am @Somy .

sammixboo (sammixboo):

~(*O*)~ Somy would be a great Ambie!!! Oh, and congrats on becoming olive ;o You will be emerald pretty soon ~(=u=~)

OpenStudy (luigi0210):

If you can take over the OS website, you can be whatever you want to be ;) And yesh, but don't cheat to green.. unlike some *cough* .. hi Sammi :D But meh, ambie, hambie, damie, cammie, have some candy :D ⓐ

OpenStudy (somy):

thnx @sammixboo lol i aim for 100 xD though....it'll take ages lmao

OpenStudy (anonymous):

my vote for @PRAETORIAN.10

OpenStudy (chmvijay):

awww Sorry i forgot @Abhisar :) i equally recommend his name to for the ambassador :)

OpenStudy (abhisar):

XD XD It's ok @chmvijay & Thnx \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (ikram002p):

@Somy <3

OpenStudy (luigi0210):

@chmvijay for ambie :3

OpenStudy (adrynicoleb):

Nuhhh I would be a terrible mod Lui xD And Vijay is a good mod xD

OpenStudy (chmvijay):

@Luigi0210 i recommend you too :) o my bad i am forgetting to many good people here :)

OpenStudy (luigi0210):

I got fired :D

OpenStudy (adrynicoleb):

They should rehire ya. @Luigi0210

OpenStudy (hari5719):

somy will be a very good ambassador indeed